OpenStudy (anonymous):

HELPP Use substitution y = 2x-5 5x-4y=14

OpenStudy (anonymous):

put y = 2x-5 in the second equation to proceed

OpenStudy (anonymous):

....

OpenStudy (anonymous):

5x-4*(2x-5)=14 can you solve now?

OpenStudy (triciaal):

replace y in the 2nd eqn with 2x+5 5x -4(2x-5)= 14 5x -8x + 20 = 14 -3x = 14-20 x = -6/-3 =2 you have x substitute in first y = 2(2) -5 = -1

OpenStudy (anonymous):

x= 2